### v3

@@ 2.1 Multi-Robot Systems
A MRS consists of multiple robots that collaborate to complete specific tasks. Unlike single-robot systems, MRS leverages the combined capabilities of multiple robots to perform complex tasks more efficiently, reliably, and flexibly [5, 75, 114]. These systems are commonly employed in applications such as search and rescue [8, 61, 44, 73], target tracking [111, 113, 106, 51, 54], environmental monitoring [78, 28], coverage and exploration [10, 81, 80, 58, 11], and warehouse automation [4, 98], where the task’s scale or complexity exceeds a single robot’s capabilities. When all robots in the team are identical and share the same functionality, the team is called a homogeneous multi-robot team. In contrast, a heterogeneous multi-robot team consists of different types of robots [71, 79, 11]. The advantages of MRS include enhanced scalability, as tasks can be distributed among robots [55, 110, 13], and increased resilience, as the failure of one robot can often be mitigated by the others [116, 112, 74, 56, 66, 115, 117, 109, 82, 50]. In contrast to designing a single, highly versatile robot, MRS usually relies on more uncomplicated, task-specific robots, reducing the cost and complexity of individual units while benefiting from collective intelligence [39]. However, these systems also present unique challenges, particularly in communication, coordination, and decision-making, as robots must operate cohesively in dynamic and uncertain environments [73]. Two primary control paradigms are commonly employed to manage the interaction and task distribution within an MRS: centralized and decentralized controllers [100, 16]. In a centralized controller, a single controller receives all the information and directs the actions of all robots in the system, allowing for optimized coordination and global planning. However, centralized systems can become a bottleneck when the group size increases and are vulnerable to single points of failure [60]. On the other hand, a decentralized controller distributes decision-making among the roots, enabling the robots to operate resiliently [75]. This approach enhances scalability and resilience but often introduces additional complexity to ensure seamless communication and coordination between robots. The choice between centralized and decentralized control depends on the specific application requirements, environmental conditions, and the desired balance between efficiency and robustness [100].
@@ 2.2.2 Multimodal LLMs
Traditional LLMs excel at processing and generating text but fall short in scenarios where understanding multiple data types is essential. Recent progress on multimodal LLMs addresses this limitation by incorporating diverse modalities, enabling them to combine textual inputs with visual, auditory, or other sensory data [102]. These models align information from different modalities into a shared semantic space, allowing for seamless integration and contextual understanding. For instance, a multimodal LLM can process visual data from a robot’s camera alongside textual commands to identify objects, navigate environments, or perform complex tasks [42, 92]. This ability to synthesize information across modalities significantly enhances their applicability, especially in robotics, where real-world interactions demand integrating various data types. By leveraging multimodal capabilities, these models push the boundaries of what LLMs can achieve, offering a new level of flexibility and adaptability.
@@ 2.3 Related Survey Papers
None of these surveys address the challenges and opportunities of integrating LLMs into MRS. While multi-agent systems provide a generalized framework for understanding roles and interactions, they are often abstract and virtual, lacking the physical embodiment and real-world constraints that characterize MRS [30]. MRS requires actual physical robots to collectively perceive, decide, and act within dynamic and uncertain environments, posing unique challenges in communication, coordination, and decision-making that go beyond the scope of virtual agents [90]. Moreover, MRS uniquely benefits from improved scalability, failure resilience, and cost-effective collective operations, making them fundamentally different from single-robot systems or general multi-agent frameworks. This gap highlights the need for a dedicated survey that explores how LLMs can facilitate communication, coordination, and collaborative task execution in MRS, providing critical insights into this emerging and impactful area of research.
@@ 4.1 High-Level Task Allocation and Planning
In addition, Brienza et al. [9] applied VLM and LLM to generate actionable plans for the robotic soccer team. Their approach involved providing the VLM coach with a training set comprising video frames paired with corresponding textual prompts, detailing tasks and constraints. The VLM coach generated schematic descriptions of the video frames along with high-level natural language plans. Two distinct LLMs further refined and synchronized these high-level plans to produce executable strategies suitable for various scenarios. In practical applications, the system selected pre-collected plans based on their similarity to the real-world situation. Additionally, RAG minimizes prompt size and mitigates hallucination, ensuring more reliable outputs. Moreover, Lykov et al. [62] developed an MRS to collect and sort colored object sets and count spherical objects. Their approach utilized a fine-tuned LLM to generate Behavior Trees (BTs) for robots to execute tasks and provide feedback to human operators regarding their behaviors. They implemented a single LLM with two LoRA adapters, each handling specific functionalities to enhance efficiency and resource compactness. In addition, Ahn et al. [2] introduced an MRS framework featuring a recovery mechanism. The LLM controller received natural language instructions and a library of low-level robot skills to generate plans for task execution. A key innovation in their system was detecting deviations from the expected task progression and performing error recovery by replanning or seeking assistance from other robots or human operators. The remaining studies in this domain can be further categorized into two key areas: multi-robot multi-task coordination and complex task decomposition, highlighting the breadth of LLM applications in MRS.
4.1.1 Multi-Robot Multi-Task
@@
Lakhnati et al. [45] proposed a framework where three heterogeneous robots aim to accomplish complex tasks instructed by human operators in VR simulation. First, each robot LLM is given an initial prompt to clarify its role and abilities. A central controller LLM analyzes human descriptions of the task and distributes them to the respective robots. Instructions from human operators can either directly specify what each robot should do (e.g., “Jupiter needs to move to the dumbbell and pick it up, Neptune and Pluto have to move to the fridge.”) or describe the tasks without assigning to specific robots (e.g., “Three dinner plates have to be put into the trash, and all agents need to end up next to the garbage bin.”). Following this line, Chen et al. [12] proposed a centralized framework where an LLM controller distributes the human instructions to a multi-robot team. They aim to make a heterogeneous multi-robot team accomplish multiple heterogeneous household tasks. However, the task distribution process they introduced is in the form of a discussion between the “Central Planner” LLM and the robot-dedicated agent LLM on each robot. The original task information is a geometric representation from a SLAM system. It is constructed into a scene context to prompt LLM. The “Central Planner” LLM first assigns each task to each robot according to its analysis. Then, each robot-dedicated agent LLM provides feedback according to the assigned task, and its robot resume is generated from the robot’s URDF code by the robot-dedicated agent LLM. If the task does not match the robot’s resume, it prompts the “Central Planner” for a reassignment. This discussion between LLMs continues until no reassignments are required. Chen et al. [15] took a step further to investigate the scalability of an LLM-based heterogeneous multi-task planning system. The efficiency and accuracy of four different communication architectures are compared, as shown in Fig. 3 under four distinct environments, including BoxNet, warehouse, and BoxLift. The results demonstrate that the HMAS-2 structure achieves the highest success rate while CMAS is the most token-efficient. On the other hand, Gupte et al. [31] proposed an LLM-based framework to solve Initial Task Allocation for a multi-robot multi-human system. In this centralized framework, LLM first generates prescriptive rules for each user’s objective and then generates experiences based on those rules for each objective. After acquiring a practical knowledge of the rules generated, the LLM’s performance is evaluated by inferencing, where the user provides instructions, and the LLM allocates the task according to the rules and experiences. Two distinct RAG workflows are leveraged in the inferencing stage to use the acquired knowledge fully. Moreover, Huang et al. [35] tested the ability of LLMs to solve the multi-robot Traveling Salesman Problem (TSP). By providing appropriate prompts, the LLM plans the optimal paths for multiple robots and generates Python code to control their movements. The study set up three frameworks: single attempt, self-debugging, where the LLM checks whether the generated Python code can be executed, and self-debugging with self-verification, where the LLM checks code executability and verifies whether the execution produces correct results. Their work reveals that LLMs perform poorly in handling such problems, with higher success rates that can only be observed in specific cases, such as the min-max multi-robot TSP.
@@ 4.1.2 Complex Task Decomposition
Xu et al. [99] proposed a two-step framework that leverages LLMs to translate complex natural language instructions into a hierarchical linear temporal logic (LTL) representation for MRS. In the first step, the LLM decomposes the instruction into a hierarchical task tree, capturing logical and temporal dependencies between subtasks to avoid errors in sequence. In the second step, a fine-tuned LLM translates each subtask into flat LTL formulas, enabling precise execution using off-the-shelf planners. This framework emphasizes the importance of temporal reasoning in decomposing complex instructions, ensuring accurate task allocation and execution for long-horizon and interdependent multi-robot tasks. In contrast to the abovementioned approaches, Obata et al. [70] adopted a slightly different approach and proposed LiP-LLM, a framework integrating LLMs with linear programming for multi-robot task planning. Instead of providing end-to-end task allocation and execution, LiP-LLM utilizes LLMs to generate a skill set and a dependency graph that maps relationships and sequential constraints among tasks. The task allocation is then solved using linear programming to optimize task distribution among robots. This hybrid approach enhances task execution efficiency and success rates by combining LLMs’ interpretative capabilities with the precision of optimization techniques. The results demonstrate the potential of integrating LLMs with traditional optimization techniques to improve the performance and coordination of MRS. On the other hand, Liu et al. [57] proposed the COHERENT framework, which utilizes a Proposal-Execution-Feedback-Adjustment (PEFA) mechanism for task planning in heterogeneous MRS. The PEFA process involves a centralized task assigner LLM that decomposes high-level human instructions into subgoals and assigns them to individual robots. Each robot evaluates the assigned subgoal, determines its feasibility, and provides feedback to the task assigner, enabling dynamic adjustments and iterative refinements in the task plan. This process bears similarities to the robot discussion mechanism in the EMOS framework proposed by Chen et al. [12], where task decomposition and assignment leverage embodiment-aware reasoning based on robot resumes. However, COHERENT emphasizes a real-time, feedback-driven approach to handle task allocation and execution, making it particularly suited for dynamic and complex multi-robot environments. Differently, Mandi et al. [65] proposed RoCo, a decentralized communication architecture for multi-robot collaboration, focusing on both high-level task planning and low-level motion planning. In the RoCo framework, each robot is equipped with an LLM that engages in dialogue with other robots to discuss and refine task strategies. This dialogue process results in a proposed sub-task plan, which is validated by the environment for feasibility. If the plan fails (e.g., due to collisions or invalid configurations), feedback is incorporated into subsequent dialogues to improve the plan iteratively. Once validated, the sub-task plan generates goal configurations for robot arms, with a centralized motion planner computing collision-free trajectories. RoCo emphasizes flexibility and adaptability in multi-robot collaboration and has been evaluated using the RoCoBench benchmark, demonstrating robust performance across diverse task scenarios. This approach highlights the synergy between decentralized LLM-driven reasoning and centralized motion planning for complex, dynamic environments.

### v4

@@ Abstract
[…]
Competing Interests: The authors declare no competing interests.
@@ 2.1 Multi-Robot Systems
A MRS consists of multiple robots that collaborate to complete specific tasks. Unlike single-robot systems, MRS leverages the combined capabilities of multiple robots to perform complex tasks more efficiently, reliably, and flexibly [5, 98, 147]. These systems are commonly employed in applications such as search and rescue [8, 78, 55, 95], target tracking [144, 146, 136, 63, 71], environmental monitoring [101, 33], coverage and exploration [12, 105, 103, 75, 13], and warehouse automation [4, 125], where the task’s scale or complexity exceeds a single robot’s capabilities. When all robots in the team are identical and share the same functionality, the team is called a homogeneous multi-robot team. In contrast, a heterogeneous multi-robot team consists of different types of robots [92, 102, 13]. The advantages of MRS include enhanced scalability, as tasks can be distributed among robots [72, 143, 16], and increased resilience, as the failure of one robot can often be mitigated by the others [149, 145, 97, 73, 85, 148, 150, 142, 106, 62, 60]. In contrast to designing a single, highly versatile robot, MRS usually relies on more uncomplicated, task-specific robots, reducing the cost and complexity of individual units while benefiting from collective intelligence [49]. However, these systems also present unique challenges, particularly in communication, coordination, and decision-making, as robots must operate cohesively in dynamic and uncertain environments [95]. Two primary control paradigms are commonly employed to manage the interaction and task distribution within an MRS: centralized and decentralized controllers [130, 20]. In a centralized controller, a single controller receives all the information and directs the actions of all robots in the system, allowing for optimized coordination and global planning. However, centralized systems can become a bottleneck when the group size increases and are vulnerable to single points of failure [77]. On the other hand, a decentralized controller distributes decision-making among the roots, enabling the robots to operate resiliently [98]. This approach enhances scalability and resilience but often introduces additional complexity to ensure seamless communication and coordination between robots. The choice between centralized and decentralized control depends on the specific application requirements, environmental conditions, and the desired balance between efficiency and robustness [130].
@@ 2.2.2 Multimodal LLMs
Traditional LLMs process only text, limiting their ability to directly interpret sensory data. In multi-robot systems, however, perception is often the primary source of situational awareness, coming from heterogeneous sensors such as onboard cameras, LiDAR, or aerial imagery. Multimodal LLMs address this gap by integrating inputs such as images, video, audio, or structured sensor data into a shared semantic space alongside language [132]. This allows robots to ground natural language reasoning in real-world observations—for example, identifying mission-relevant objects, understanding spatial relationships, or updating plans based on live visual feedback [53, 117].
Recent advances have led to the emergence of Vision-Language Models (VLMs), which couple perception and reasoning, and Vision-Language-Action Models (VLAs), which go further by linking perception and reasoning directly to executable actions. These models extend the role of LLMs in MRS from purely symbolic reasoning to perception-grounded decision-making and, in the case of VLAs, to closed-loop perception–action execution [141]. As discussed in Sec. 6.1, this evolution opens new possibilities for coordinated, perception-driven, and time-critical decision-making in multi-robot teams.
@@ 2.3 Related Survey Papers
None of these surveys addresses the challenges and opportunities of integrating LLMs into MRS. While multi-agent systems provide a generalized framework for understanding roles and interactions, they are often abstract and virtual, lacking the physical embodiment and real-world constraints that characterize MRS [35]. MRS requires actual physical robots to collectively perceive, decide, and act within dynamic and uncertain environments, posing unique challenges in communication, coordination, and decision-making that go beyond the scope of virtual agents [114]. Moreover, MRS uniquely benefits from improved scalability, failure resilience, and cost-effective collective operations, making them fundamentally different from single-robot systems or general multi-agent frameworks. This gap highlights the need for a dedicated survey that explores how LLMs can facilitate communication, coordination, and collaborative task execution in MRS, providing critical insights into this emerging and impactful area of research.
@@ 4.1 High-Level Task Allocation and Planning
In addition, Brienza et al. [10] applied VLM and LLM to generate actionable plans for the robotic soccer team. Their approach involved providing the VLM coach with a training set comprising video frames paired with corresponding textual prompts, detailing tasks and constraints. The VLM coach generated schematic descriptions of the video frames along with high-level natural language plans. Two distinct LLMs further refined and synchronized these high-level plans to produce executable strategies suitable for various scenarios. In practical applications, the system selected pre-collected plans based on their similarity to the real-world situation. Additionally, RAG minimizes prompt size and mitigates hallucination, ensuring more reliable outputs. Moreover, Lykov et al. [79] developed an MRS to collect and sort colored object sets and count spherical objects. Their approach utilized a fine-tuned LLM to generate Behavior Trees (BTs) for robots to execute tasks and provide feedback to human operators regarding their behaviors. They implemented a single LLM with two LoRA adapters, each handling specific functionalities to enhance efficiency and resource compactness. In addition, Ahn et al. [2] introduced an MRS framework featuring a recovery mechanism. The LLM controller received natural language instructions and a library of low-level robot skills to generate plans for task execution. A key innovation in their system was detecting deviations from the expected task progression and performing error recovery by replanning or seeking assistance from other robots or human operators. Li and Zhou [64] propose LLM-Flock, an approach that integrates decentralized LLM-driven planning with an influence-based consensus protocol for multi-robot formation control. This method mitigates inconsistent plan and behavior by refining local plans iteratively based on the influence of neighbors, achieving stable and adaptable flocking. Validated through simulations and drone experiments, their approach significantly enhances the coherence of multi-robot formations. The remaining studies in this domain can be further categorized into two key areas: multi-robot multi-task coordination and complex task decomposition, highlighting the breadth of LLM applications in MRS.
4.1.1 Multi-Robot Multi-Task Planning
@@
Lakhnati et al. [56] proposed a framework where three heterogeneous robots aim to accomplish complex tasks instructed by human operators in VR simulation. First, each robot LLM is given an initial prompt to clarify its role and abilities. A central controller LLM analyzes human descriptions of the task and distributes them to the respective robots. Instructions from human operators can either directly specify what each robot should do (e.g., “Jupiter needs to move to the dumbbell and pick it up, Neptune and Pluto have to move to the fridge.”) or describe the tasks without assigning to specific robots (e.g., “Three dinner plates have to be put into the trash, and all agents need to end up next to the garbage bin.”). Following this line, Chen et al. [15] proposed a centralized framework where an LLM controller distributes the human instructions to a multi-robot team. They aim to make a heterogeneous multi-robot team accomplish multiple heterogeneous household tasks. However, the task distribution process they introduced is in the form of a discussion between the “Central Planner” LLM and the robot-dedicated agent LLM on each robot. The original task information is a geometric representation from a SLAM system. It is constructed into a scene context to prompt LLM. The “Central Planner” LLM first assigns each task to each robot according to its analysis. Then, each robot-dedicated agent LLM provides feedback according to the assigned task, and its robot resume is generated from the robot’s URDF code by the robot-dedicated agent LLM. If the task does not match the robot’s resume, it prompts the “Central Planner” for a reassignment. This discussion between LLMs continues until no reassignments are required. Lim et al. [69] presented a centralized LLM-enabled control framework for dynamic task assignment in multi-robot manufacturing systems. The architecture consists of three main components. The Central Controller Agent (CCA) acts as a high-level decision-maker, leveraging LLM reasoning to reassign tasks when disruptions occur. The Robot Agents execute tasks using configuration-defined capabilities and report failures to the CCA. The Sensor Module provides real-time perceptual input to ensure context-aware decisions. When a robot failure is detected, the CCA formulates a structured prompt containing system constraints and objectives, and sends it to the LLM to generate a valid new configuration for an alternative robot. If the LLM-generated configuration violates task constraints, structured feedback is appended, and the process is repeated until success is achieved. This loop ensures that reassignment is both feasible and safe. Once a valid configuration is found, the task is reallocated, and execution resumes seamlessly. Moreover, Jiang et al. [48] introduced a decentralized swarm robotics framework that leverages LLMs to enable spontaneous inter-robot communication and coordination. Each robot maintains its own LLM session and utilizes natural language to discover peers, share information, and coordinate collaborative actions without predefined roles. The system demonstrates that LLM-driven communication can give rise to emergent social behaviors such as negotiation, assistance, and conflict resolution, pushing swarm robotics toward more human-like, adaptive coordination. Chen et al. [18] took a step further to investigate the scalability of an LLM-based heterogeneous multi-task planning system. The efficiency and accuracy of four different communication architectures are compared, as shown in Fig. 3 under four distinct environments, including BoxNet, warehouse, and BoxLift. The results demonstrate that the HMAS-2 structure achieves the highest success rate while CMAS is the most token-efficient. On the other hand, Gupte et al. [37] proposed an LLM-based framework to solve Initial Task Allocation for a multi-robot multi-human system. In this centralized framework, LLM first generates prescriptive rules for each user’s objective and then generates experiences based on those rules for each objective. After acquiring a practical knowledge of the rules generated, the LLM’s performance is evaluated by inferencing, where the user provides instructions, and the LLM allocates the task according to the rules and experiences. Two distinct RAG workflows are leveraged in the inferencing stage to use the acquired knowledge fully. Moreover, Huang et al. [41] tested the ability of LLMs to solve the multi-robot Traveling Salesman Problem (TSP). By providing appropriate prompts, the LLM plans the optimal paths for multiple robots and generates Python code to control their movements. The study set up three frameworks: single attempt, self-debugging, where the LLM checks whether the generated Python code can be executed, and self-debugging with self-verification, where the LLM checks code executability and verifies whether the execution produces correct results. Their work reveals that LLMs perform poorly in handling such problems, with higher success rates that can only be observed in specific cases, such as the min-max multi-robot TSP. Chen et al. [14] presented a multi-agent LLM-based framework for robotic autonomy, integrating task analysis, robot design, and RL path planning. Experiments demonstrate the effectiveness of robot configuration generation and the feasibility of control strategies, showcasing significant improvements in the scalability and flexibility of LLM-driven robotic system development. Wan et al. [115] introduced a supervised fine-tuned LLM-based planning framework using the novel MultiPlan dataset for embodied multi-robot collaboration. Extensive indoor and outdoor field experiments illustrate significant improvements in task decomposition, robustness, and adaptability to novel environments compared to baseline LLM planners.
@@ 4.1.2 Complex Task Decomposition
Xu et al. [127] proposed a two-step framework that leverages LLMs to translate complex natural language instructions into a hierarchical linear temporal logic (LTL) representation for MRS. In the first step, the LLM decomposes the instruction into a hierarchical task tree, capturing logical and temporal dependencies between subtasks to avoid errors in sequence. In the second step, a fine-tuned LLM translates each subtask into flat LTL formulas, enabling precise execution using off-the-shelf planners. This framework emphasizes the importance of temporal reasoning in decomposing complex instructions, ensuring accurate task allocation and execution for long-horizon and interdependent multi-robot tasks. In contrast to the abovementioned approaches, Obata et al. [90] adopted a slightly different approach and proposed LiP-LLM, a framework integrating LLMs with linear programming for multi-robot task planning. Instead of providing end-to-end task allocation and execution, LiP-LLM utilizes LLMs to generate a skill set and a dependency graph that maps relationships and sequential constraints among tasks. The task allocation is then solved using linear programming to optimize task distribution among robots. This hybrid approach enhances task execution efficiency and success rates by combining LLMs’ interpretative capabilities with the precision of optimization techniques. The results demonstrate the potential of integrating LLMs with traditional optimization techniques to improve the performance and coordination of MRS. On the other hand, Liu et al. [74] proposed the COHERENT framework, which utilizes a Proposal-Execution-Feedback-Adjustment (PEFA) mechanism for task planning in heterogeneous MRS. The PEFA process involves a centralized task assigner LLM that decomposes high-level human instructions into subgoals and assigns them to individual robots. Each robot evaluates the assigned subgoal, determines its feasibility, and provides feedback to the task assigner, enabling dynamic adjustments and iterative refinements in the task plan. This process bears similarities to the robot discussion mechanism in the EMOS framework proposed by Chen et al. [15], where task decomposition and assignment leverage embodiment-aware reasoning based on robot resumes. However, COHERENT emphasizes a real-time, feedback-driven approach to handle task allocation and execution, making it particularly suited for dynamic and complex multi-robot environments. Huang et al. [42] proposed LAN2CB (Language to Collective Behavior), a framework that converts natural language mission descriptions into executable multi-robot code. The system performs mission parsing, dependency analysis, and behavior tree construction, then generates task code using a library of goal generation, allocation, and motion primitives. LAN2CB supports dynamic replanning by updating the behavior tree and regenerating code when triggers occur, enabling robust adaptation in changing environments. Experiments on nine diverse scenarios in both simulation and real-world tests demonstrated high success rates and resilience to complex dependencies and environmental events.
Liang et al. [68] proposed a communication-based, feedback-driven framework that enhances multi-robot cooperation through a novel retrospective actor-critic paradigm. Their system comprises two large language models: $\text{LLM}_{1}$, which facilitates real-time collaborative discussion among robots to devise initial plans, and $\text{LLM}_{2}$, which serves as both an action critic and proposer. After executing a task or simulation, $\text{LLM}_{2}$ conducts a retrospective evaluation based on environmental feedback, critiques the earlier plans, and proposes improved alternatives. The framework leverages short-term and long-term memory to retain only the most recent retrospectives, thus maintaining contextual relevance while minimizing computational overhead. This approach demonstrates improved performance on RoCoBench tasks compared to baselines, highlighting the value of integrating LLM-based reflection into the decision-making pipeline for dynamic and uncertain multi-robot environments. Differently, Mandi et al. [84] proposed RoCo, a decentralized communication architecture for multi-robot collaboration, focusing on both high-level task planning and low-level motion planning. In the RoCo framework, each robot is equipped with an LLM that engages in dialogue with other robots to discuss and refine task strategies. This dialogue process results in a proposed sub-task plan, which is validated by the environment for feasibility. If the plan fails (e.g., due to collisions or invalid configurations), feedback is incorporated into subsequent dialogues to improve the plan iteratively. Once validated, the sub-task plan generates goal configurations for robot arms, with a centralized motion planner computing collision-free trajectories. RoCo emphasizes flexibility and adaptability in multi-robot collaboration and has been evaluated using the RoCoBench benchmark, demonstrating robust performance across diverse task scenarios. This approach highlights the synergy between decentralized LLM-driven reasoning and centralized motion planning for complex, dynamic environments. Peng et al. [93] introduced a knowledge-augmented automated MILP formulation framework for multi-robot task allocation and scheduling. Their approach leverages locally deployed LLMs, combined with a structured knowledge base, to extract spatiotemporal constraints from natural language and translate them into executable MILP code. The two-step pipeline first interprets fuzzy natural instructions into mathematical constraints and then generates optimized Gurobi-compatible code. Notably, the entire process is locally deployable, ensuring data privacy in sensitive industrial settings. Experimental results on aircraft skin manufacturing tasks demonstrate that the framework achieves high accuracy and efficiency in both model construction and code generation, highlighting the potential of integrating LLMs with symbolic optimization techniques for practical, privacy-sensitive MRS applications. Cladera et al. [19] demonstrated air-ground collaborative planning using LLM-based semantic reasoning and mapping for language-specified missions. Their decentralized system enables dynamic mission adaptation, as validated through extensive field experiments in both urban and rural settings, highlighting effective semantic-based collaboration and robust task execution. Gupta et al. [36] proposed a hierarchical tree-based mission planning approach that leverages LLMs for automated decomposition into manageable subtasks, explicitly tailored for heterogeneous robot teams. The proposed heuristic method efficiently schedules tasks, considering robot-specific constraints, and demonstrates flexibility and scalability across varied, complex missions.